OpenStudy (anonymous):

Check my answer? Solve. 2x2 – 8x – 16 = 0 x = 2 ± 2 sqrt 3 x = 4 ± 2 sqrt 5 x = 2 ± 4 sqrt 2 x = 4 ± 2 sqrt 3 I think its the thrid one!

OpenStudy (anonymous):

*third

OpenStudy (anonymous):

unfortunately no, and i bet i know exactly the mistake you made, it is easy to do

OpenStudy (anonymous):

i bet you divided by 2 in the numerator right?

OpenStudy (anonymous):

lol yea i did.

OpenStudy (anonymous):

unfortunately you cannot divide inside the radical by 2, you have to write in simplest radical form first lets do it the easy way

OpenStudy (anonymous):

you can check these on you graphing calculator if you have one

hero (hero):

Hint: Factor by grouping

OpenStudy (anonymous):

@Outkast3r09 i dont have one:/

OpenStudy (anonymous):

@Hero the answer is a radican

OpenStudy (precal):

quadratic formula would also work

OpenStudy (anonymous):

first divide everything by 2 to get \[x^2-4x-8=0\] the we can use the quadratic formula or what is easier is to complete the square \[x^2-4x-8=0\] \[x^2-4x=8\] \[(x-2)^2=8+4=12\] \[x-2=\pm\sqrt{12}=\pm2\sqrt{3}\] \[x=2\pm2\sqrt{3}\]

Parth (parthkohli):

No... no factoring... QUADRATIC FORMULA.

OpenStudy (anonymous):

now lets use the formula and see what happens to get the same answer the formula is harder because it forces a denominator on you, but there is none

OpenStudy (anonymous):

okay thanks

OpenStudy (precal):

either method would work, you would have to simplify at the end

OpenStudy (anonymous):

\[x^2-4x-8=0\] \[a=1,b=-4,c=-8\] \[x=\frac{4\pm\sqrt{(-4)^2-4\times 1\times -8}}{2}\] \[x=\frac{4\pm\sqrt{48}}{2}\] now you cannot divide 48 by 2 to get \(\sqrt{24}\) because it is not \(48\) is is \(\sqrt{48}\) so you have to write it in simplest radical form before you can divide

hero (hero):

@outkast3r09, that did not stop me from factoring by grouping

OpenStudy (anonymous):

\[\sqrt{48}=4\sqrt{3}\]so now you have \[\frac{x=4\pm4\sqrt{3}}{2}\]and NOW you can divide both terms by 2 to get \[x=2\pm2\sqrt{3}\]

OpenStudy (precal):